OpenStudy (anonymous):

Solve the quadratic equation. 48 - 8x2 = 0 A. ± √6 B. ± 2√2 C. ± (2√2)/√2 D. ± (4√3)/(2√2)

Directrix (directrix):

48 - 8x2 = 0 Divide each member of the equation by 8 6 - x² = 0 - x² + 6 = 0 Add (-6) to both sides 6 - x² - 6 = 0 - 6 - x² = -6 Multiply each side by -1 x² = 6 x = ± √6